OpenStudy (anonymous):

Evaluate: 64^3/4, give exact answer, show work please

OpenStudy (anonymous):

65536

OpenStudy (anonymous):

64^3/4 = (2^4)^3/4 =2^3 =8

OpenStudy (anonymous):

let me rephrase that. 63^(3/4)

OpenStudy (anonymous):

sorry still 64 not 63